I really need help setting this problem up. I can't figure it out!
Problem:
The perimeter of a rectangle is 20 feet. The width is 4 feet less than the length. Find the dimensions of the rectangle. Then find the area.

Answer by smaraya(8)

let length be x
Let width be x-4
P=perimeter
P=2(l+b)
20=2(x+x-4)
20=2x+2x-8
20+8=4x
x=7
Length is 7 ft
Width is 3 ft